A 200-g Nerf dart is shot from a dart gun with a kinetic energy of 3.60 J. What was the initial of the speed dart?
1
Expert's answer
2020-12-21T11:02:52-0500

By definition, the kinetic energy is



"K = \dfrac{mv^2}{2}"

where "m = 200g = 0.2kg" is the mass of the dart, and "v" is its speed. Since "K = 3.6J", obtain the following speed:


"v = \sqrt{\dfrac{2K}{m}}\\
v = \sqrt{\dfrac{2\cdot 3.6}{0.2}} = 6m/s"

Answer. 6m/s.